OpenStudy (anonymous):

Liquid A has 1rams of element X per liter. Liquid B has 1 milligrams of element Y per cm^3 (there are 10milligrams in a gram and 10 cm^3 in a liter). It is desired to make 1grams of molecule Z.

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

ok

OpenStudy (perl):

damn, draw button doesnt work

OpenStudy (perl):

can you tag me again, brb

OpenStudy (perl):

im back

OpenStudy (anonymous):

.2

OpenStudy (perl):

oh wait, we want 120 grams of z

OpenStudy (anonymous):

oh right right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i can see that

OpenStudy (perl):

an easier way: 2/18 is the fractional amount of X in Z , and 16/18 is the fraction of the amount of Y in Z. this is true always. (if you multiply by 18, you get 2 and 16 respectively) now if you have 120 grams, this fractional amount wont change 120 * 2/18 = grams of X in Z 120 * 16/18 = grams of Y in Z

OpenStudy (perl):

120 * 2/18 = 13.333... grams of X 120 * 16/18 = 106.666... grams of Y

OpenStudy (perl):

ok now we can go back and use our conversion factor to find how many liters of liquid A and B we need

OpenStudy (anonymous):

alright

OpenStudy (perl):